OpenStudy (anonymous):

1 over 3(c-2)= 7 over 3

jhonyy9 (jhonyy9):

and solve this for c ?

OpenStudy (anonymous):

yeah

jhonyy9 (jhonyy9):

so there are 1 7 ------ = --- 3(c-2) 3 than cross multiplie what will get ?

jhonyy9 (jhonyy9):

do you know it ?

OpenStudy (anonymous):

i dont understand how to do it

jhonyy9 (jhonyy9):

please this is easy cross multiplie and will get 3=21(c-2) so can you continue it ?

OpenStudy (anonymous):

so then i mulitpile

jhonyy9 (jhonyy9):

divide both sides by 3 and will get 1=7(c-2) make the calcules and will get 1=7c -14 add to both sides 14 1+14=7c-14+14 15=7c

jhonyy9 (jhonyy9):

and how many will be c ?

OpenStudy (anonymous):

let me work it out

OpenStudy (anonymous):

this is my first year doing alegbra

jhonyy9 (jhonyy9):

ok i will wait your reply

OpenStudy (anonymous):

thnks

jhonyy9 (jhonyy9):

so than 15=7c how you need to continue it ?

OpenStudy (anonymous):

2.1 correct

jhonyy9 (jhonyy9):

write it please how you get it

OpenStudy (anonymous):

15 divided by 7 which equals 2.14 then round it equal 2.1

jhonyy9 (jhonyy9):

yes is right but if not is necessary calcule the fraction than you can write it just like c= 15/7 ok ?

OpenStudy (anonymous):

thnk you so much

jhonyy9 (jhonyy9):

was my pleasure good luck bye